OpenStudy (anonymous):

divide x^2+8x+15 _______________divided x^2+6x+9/4x^2-5x-6 4x^2+23x+15

OpenStudy (rogue):

Reciprocate the second fraction and multiply the two fractions. Try to factor things out to make it simpler.

OpenStudy (anonymous):

(x-2)^2(4x+3)/(x+3)

OpenStudy (anonymous):

Factored mostly with ac grouping except x^2+6x+9 which is simply (x+a)^2

OpenStudy (rogue):

\[\frac {x ^{2} + 8x +25}{4x ^{2} + 23x +15} \div \frac {x^{2} + 6x + 9}{4x^{2} - 5x -6}\]

OpenStudy (anonymous):

1st denominator factored by grouping like so: (4x^2 + 20x + 3x + 15) = 4x(x+5)+3(x+5)=(x+5)(4x+3)

OpenStudy (rogue):

\[\frac {x ^{2} + 8x +25}{4x ^{2} + 23x +15} * \frac {4x^{2} - 5x -6}{x^{2} + 6x + 9}\]

OpenStudy (rogue):

And then factor & simplify as Rick says.

OpenStudy (anonymous):

2nd denominator like: 4x^2-8x+3x-6=4x(x-2)+3(x-2)=(x-2)(4x+3)

OpenStudy (rogue):

+15*

OpenStudy (anonymous):

Sorry?

OpenStudy (rogue):

I typed the first quadratic wrong, was correcting it to 15, instead of 25.

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

Then its just a matter of cancelling out like factors